When I first went to college, the probability that my mother was going to call me up during a given week, complaining, was 2/3. The probability of my stepfather was going to call me up during a given week, complaining, was 3/7. The probability that both were going to call me during a given week, complaining, was 2/7.

a) What was the probability that Mom was going to call, given that Pop had called?

b) What was the probability that I was going to get at least one call from a parent during a given week?

c) Are these two events independent? Why or why not?
Answer #1

a)P(Mom |Pop) =P(both)/P(pop) =(2/7)/(3/7)=2/3

b)P(At least one call) =P(Pop)+P(Mom)-P(both) =(3/7)+(2/3)-(2/7)=17/21

c) since P(Mom)*P(Pop) =(2/3)*(3/7)=2/7 =P(both),

the events are independent